40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Mon Nov 3 15:16:28 2014, Dmass, Noise Hunting, Noise Budget, New beat spectrum: ErrSigs.pngBeatNz.png
    Reply  Mon Nov 3 16:15:21 2014, Dmass, Noise Hunting, Noise Budget, Beat spectrum not ltd by  
       Reply  Mon Nov 3 20:27:28 2014, Dmass, Noise Hunting, Noise Budget, ISS servo fail? 
          Reply  Wed Nov 19 14:54:36 2014, nicolas, Noise Hunting, Noise Budget, ISS servo fail?  
       Reply  Tue Dec 2 22:51:52 2014, Dmass, Noise Hunting, Noise Budget, Beat spectrum not ltd by  PDH_demod_chain.png
          Reply  Sat Sep 12 14:25:05 2015, Dmass, Noise Hunting, Noise Budget, Beat spectrum not ltd by  
    Reply  Mon Nov 3 20:36:14 2014, Dmass, Noise Hunting, Noise Budget, New beat spectrum: 
       Reply  Tue Nov 25 22:45:43 2014, Dmass, Noise Hunting, Noise Budget, New beat spectrum: 
          Reply  Wed Dec 3 17:02:31 2014, Dmass, Noise Hunting, Noise Budget, Noise budget cleaning up 
             Reply  Thu Dec 4 02:45:49 2014, Dmass, Noise Hunting, Noise Budget, Noise budget cleaning up PDH_sensing_nz.pngLB1005inputNoise.png
                Reply  Sun Dec 7 12:42:55 2014, rana, Noise Hunting, Noise Budget, Scatter Noise hump is RAM 
Message ID: 1165     Entry time: Mon Nov 3 20:36:14 2014     In reply to: 1162     Reply to this: 1168
Author: Dmass 
Type: Noise Hunting 
Category: Noise Budget 
Subject: New beat spectrum: 

Noise floor is marconi noise in PLL (see elog:1049)

Explored anomalous intensity noise with Chris. Were able to get sizable decrease in the noise by tuning the L/2 plates on the input beam.

No change to beat spectrum level.

 

With both temperature loops engaged, beat drift rate is 400Hz/sec.

RMS frequency noise from PDH loops is ~200Hzrms - mostly from Fourier frequencies 100kHz to 500kHz

 

Might be able to get away with 2kHz Marconi range if we take care of the drift rate => this problem might disappear @ 120K because of the zero crossing

 

Would need a 400 Hz marconi range (assuming no other noise pops up) to see coating thermal noise @ 1kHz

 

What noise/range did the DFD box from the 40m end up hitting?

ELOG V3.1.3-